OpenStudy (anonymous):

You need to lift a 200 kg box 5 meters, but it is too heavy. You decide to push it up an incline that makes a 25 degree angle with the ground. With what force do you need to push the box?

OpenStudy (caominhim):

if you use energy equations calculate the potential energy at the top of the ramp and then use the work equation (w =FDcos(theta)) to find the force or you can find the gravitational component that pushes down the ramp and you have to counter that force with the same amount to bring balance to the force ;D

OpenStudy (caominhim):

this all assumes no friction

OpenStudy (shamim):

Weight of 200kg box=mg=200*9.8N. Right?

OpenStudy (shamim):

Component of weight along the inclined plane=200*9.8*sin25=?

OpenStudy (shamim):

So applied force=200*9.8*sin25=?

OpenStudy (caominhim):

yep